OpenStudy (anonymous):

medal question 1. solve the equation 3m - 4 = 26 A. M = 30 B. M = 22/3 C. M = 10 D. M = 17

OpenStudy (anonymous):

Substitute the values given for x & y in the equation. For example: 17 = 4(5)+3 17 does not equal 23 so A is not the answer

OpenStudy (anonymous):

5 = 4(17)+ 3 no way 19 = 4(4)+ 3 19 = 16 + 3 this is true so it's the solution

OpenStudy (anonymous):

Do the same with #2.

OpenStudy (anonymous):

@Whiterabbit541

OpenStudy (anonymous):

@Indivicivet

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@graysondeblock

zepdrix (zepdrix):

\[\Large\bf\sf 3m-4\quad=\quad 26\]So we want to get the m alone by itself. We'll try to perform the opposite operation on everything that's happening on the side of the equation involving the m. So see how we're subtracting 4 on the left side? If we add 4 to each side, what happens? They cancel on the left side, yes?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large\bf\sf 3m-4\color{orangered}{+4}\quad=\quad 26\color{orangered}{+4}\]

zepdrix (zepdrix):

\[\Large\bf\sf 3m\quad=\quad 30\]Ok make sense so far?

OpenStudy (anonymous):

yea

zepdrix (zepdrix):

Since the 3 is multiplying the m, what would be the inverse of multiplication? How can we get rid of that 3?

OpenStudy (anonymous):

divide

zepdrix (zepdrix):

Good good good, let's divide each side by 3.

OpenStudy (anonymous):

m = 10?

zepdrix (zepdrix):

\[\Large\bf\sf \frac{3m}{3}\quad=\quad \frac{30}{3}\]

zepdrix (zepdrix):

Yes, good job! :)